# Quantum simulating continuum field theories with large-spin lattice models

### v1

@@ Abstract
Simulating the real-time dynamics of quantum field theories (QFTs) is one of the most promising applications of quantum simulators. Regularizing a bosonic QFT for quantum simulation purposes typically involves a truncation in Hilbert space in addition to a discretization of space. Here, we discuss how to perform such a regularization of scalar QFTs using multi-level or qudit systems, and show that this enables quantitative predictions in the continuum limit by extrapolating results obtained for large-spin lattice models. With extensive matrix-product state simulations, we numerically demonstrate the sequence of extrapolations that leads to quantitative agreement of observables for the integrable sine-Gordon (sG) QFT. We further show how to prepare static and moving soliton excitations, and analyze their scattering dynamics, in agreement with a semi-classical model and analytical predictions. Finally, we illustrate how a non-integrable perturbation of the sG model gives rise to dynamics reminiscent of string breaking and plasma oscillations in gauge theories. Our methods are directly applicable in state-of-the-art analog quantum simulators, opening the door to implementing a wide variety of scalar field theories and tackling long-standing questions in non-equilibrium QFT like the fate of the false vacuum.
@@ I Introduction
Quantum field theories (QFTs) play an essential role in theoretical physics, ranging from effective low-energy descriptions in condensed matter [1] to fundamental models of nature in particle physics [2]. Understanding the physics of QFTs, especially in out-of-equilibrium scenarios, is often guided by computationally challenging numerical simulations, and therefore QFTs constitute an ideal target for quantum simulation [3]. Recent quantum simulation experiments are starting to probe the boundary of classical simulations [4, 5, 6, 7, 8], with the exciting prospect to provide genuine insights into the equilibrium and non-equilibrium properties of quantum many-body systems. However, it remains an outstanding challenge to faithfully simulate physical phenomena of QFTs, such as string-breaking in gauge theories [9, 10, 11, 12].
QFTs are typically formulated in terms of continuous variables living on continuous space-time – in contrast to, e.g., many-body spin models on spatial lattices. While Lagrangian space-time discretizations are well established in numerical lattice field theory [13], digital quantum simulations are typically based on a Hamiltonian formulation with finite-dimensional Hilbert spaces and thus require truncations of bosonic field variables [14, 15]. In this context, it appears natural to employ multi-level systems as realized with trapped ions [16] or proposed with Rydberg atoms [17, 18]. Alternatively, QFTs can also arise as low-energy effective field theories in analog quantum simulations [19, 20]. However, the question how to perform controlled truncations of bosonic field variables in an analog approach has remained largely unexplored.
@@
Figure 1: Overview. (a) $d$-dimensional square lattice of particles with nearest-neighbor interactions. We assume each particle to be described by a large spin of length $J$, with $\langle\hat{J}_{+}\rangle\sim\langle e^{i\hat{\varphi}}\rangle$ and $\langle\hat{J}_{z}\rangle\sim\langle\hat{\pi}\rangle$ identifying field theory and lattice operators, as described in Sec. II. Continuum physics is extracted by a sequence of extrapolations in spin length $J$, system size $N$ and lattice spacing $a$ [see Fig. 2]. (b) Sketch of real-time scattering dynamics from large-spin models. We show how to faithfully implement 1. the vacuum (Sec. III), 2. quasi-particle wave-packets and 3. their propagation (Sec. IV), as well as 4. collisions (Sec. V resp. VI), resulting in phase shifts resp. particle production for the example of an integrable resp. non-integrable 1D QFT.
@@
In this work, we address this question and show how to quantitatively recover the continuum limit of scalar quantum field theories for an explicit truncation of the bosonic field variables. Motivated by a recent proposal to employ large-spin Rydberg atom arrays [18], our approach is based on a generalized large-spin Heisenberg Hamiltonian, which makes it immediately applicable to analog quantum simulations. We show how appropriate extrapolations [illustrated in Fig. 1(a)] enable the accurate implementation of general scalar field theories, including variants of the paradigmatic sine-Gordon (sG) model. To demonstrate our proposal, we perform extensive benchmark simulations for the 1D sG model employing tensor-network methods. In equilibrium, we find quantitative agreement with analytical predictions for ground and excited state correlation functions in the continuum limit. Out of equilibrium, we discuss how to create solitonic excitations and find that their spreading induced by quantum fluctuations is well described by a phenomenological semi-classical model. We further simulate the non-equilibrium dynamics of scattering pairs of soliton wave-packets, in agreement with predictions from the known S-matrix. Finally, we consider a perturbed sG model, where a quantum simulator would provide access to a spatially and temporally resolved picture of scattering dynamics in regimes that potentially lie beyond the reach of traditional methods.
@@
Before delving into the details, the following Section II provides an overview of our approach, including a discussion of its broad applicability and relevance for emulating condensed matter and high-energy physics. We also provide a brief summary of our main results, together with the structure of the rest of this paper, which is also illustrated in Fig. 1.
@@ II Scalar field theories from large-spin models
defined in terms of the continuum field operators $\hat{\varphi}(x)$ and their conjugate momentum $\hat{\pi}(x)$ which fulfill the canonical commutation relation $[\hat{\varphi}(x),\hat{\pi}(y)]=i\delta(x-y)$. Different models are distinguished by the form of self-interactions as described by the potential $V$.
@@ II.1 Large-spin lattice models
In App. A, we briefly review a possible experimental implementation of the above Hamiltonian using the SO(4)-symmetric high-dimensional manifold of Rydberg atoms recently discussed in Ref. [18]. In this context the spin length $J=\frac{n-1}{2}$ can be controlled by the principal quantum number $n$ of the employed Rydberg excitation, and all coupling parameters $\Omega,\Delta,\chi,\lambda_{\kappa},V_{\text{nn}}$ can be experimentally tuned by controlling atom-light interactions or rearranging the lattice geometry. Other experimentally relevant setups that give rise to related large-spin models include dipolar molecules [21], magnetic atoms [22, 23] and hybrid digital-analog proposals using qubit clusters of Rydberg atoms [24]. More generally, large-spin models can be implemented with general qudit computers [25] such as those realized with trapped ions [16] as well as hybrid qubit-oscillator processors based on cavity QED [26].
@@ II.2.1 Large-spin limit
For a square lattice regularization of $\hat{H}_{V}$ with lattice spacing $a$, one typically considers dimensionless fields ${\hat{\varphi}}_{i}$ and ${\hat{\pi}}_{i}$, that obey $[{\hat{\varphi}}_{i},{\hat{\pi}}_{j}]=i\delta_{ij}$. As a key ingredient to represent scalar fields by spins, we instead employ compactified fields where $\hat{\varphi}_{i}\in[0,2\pi]$ is restricted periodically. This means we work with the well-defined vertex operator $e^{i\hat{\varphi}}$, which obeys $[\hat{\pi}_{i},e^{i{\hat{\varphi}}_{j}}]=\delta_{ij}e^{i{\hat{\varphi}}_{j}}$ and thus acts as a raising operator on the eigenstates |$\ket{m}_{\pi}$ of $\hat{\pi}$ (at every lattice site). These operators can be regularized on the finite-dimensional Hilbert space of a spin-$J$ by identifying [27, 28]
@@
Here the $12J+$ 1 eigenstates |$\ket{m}$ of $\hat{J}_{z}$ ($m\in[-J,\dots,J]$) provide a natural restriction $|m|\leq J$ for the momentum eigenstates {|$\{\ket{m}_{\pi}\}$ of $\hat{\pi}$ at every lattice site.
@@
$\frac{\hat{J}_{\pm}}{\sqrt{J(J+1)}}\ket{m}$ $=\frac{\sqrt{J(J+1)-m(m\pm 1)}}{\sqrt{J(J+1)}}\ket{m\pm 1}$
$=\left[1-\mathcal{O}\left(\frac{m(m\pm 1)}{J(J+1)}\right)\right]\ket{m\pm 1}$ (4)
$\longleftrightarrow e^{\pm i\hat{\varphi}}\ket{m}_{\pi}$ =|$=\ket{m\pm 1}_{\pi}\;.$
Sending $J\rightarrow\infty$ therefore produces the desired lattice field operators. Note that this scaling renders the Ising term ($\propto V_{\text{nn}}\hat{J}_{z}^{(i)}\hat{J}_{z}^{(j)}$) in $\hat{H}_{\text{latt}}$ insignificant in comparison to the “flip-flop” interactions ($\propto V_{\text{nn}}\hat{J}_{+}^{(i)}\hat{J}_{-}^{(j)}$), see App. E.1.
@@ II.2.2 Numerical extrapolations
Conceptually, the large-$J$ limit of the $d$-dimensional lattice Hamiltonian $\hat{H}_{\text{latt}}$ can be translated to a continuum model by increasing the number of lattice sites $N$ and decreasing a fictitious lattice spacing $a$, while keeping the total volume $V=L^{d}$ with $L=Na$ fixed. Afterwards one can also remove the IR regulator by sending $L\rightarrow\infty$.
@@
To obtain the results shown in Fig. 2, we start by fixing the microscopic lattice parameters of $\hat{H}_{\text{latt}}$ at finite system size $N$, and compute the expectation value of the relevant observables for varying $J$. The numerical results are then fitted with a linear function in $1/J(J+1)$ [see Fig. 2(a)], from which we retrieve the results in the asymptotic limit $1/J(J+1)\rightarrow 0$, equivalent to the large spin-length limit $J\rightarrow\infty$ [see App. C for further details]. In a second step, the asymptotic values obtained for different $N$ are fitted with a quadratic function in $1/N$ [see Fig. 2(b)]. From this additional fit we retrieve the thermodynamic limit $1/N\rightarrow 0$, equivalent to $N\rightarrow\infty$. For small enough lattice spacing $a\rightarrow 0$, we expect scaling behaviour indicative of the continuum limit, which can be compared to theoretical expectations.
@@
Note that we do not tune the lattice spacing $a$ as part of the physical geometry, but we instead vary a dimensionless control parameter $aM_{0}$ – which is a function of the microscopic parameters as discussed below. This parameter sets the mass scale $M_{0}$ of the simulated QFT at the UV cutoff $1/a$ determined by our microscopic lattice model. When all extrapolations are done correctly for sufficiently small values of $aM_{0}$, the system develops a large correlation length $\xi\sim 1/{M_{0}}\gg a$, such that the extrapolated results become insensitive to the lattice regularization, and indeed agree with the continuum predictions.
@@
Figure 2: Extrapolating to the continuum limit of the simulated sine-Gordon model. (a)-(c) Extrapolation to the continuum value of the mass gap $\Delta E$. We consider the $d=1$ lattice Hamiltonian $\hat{H}_{\text{latt}}$ Eq. (II.1) and determine the mass gap $\Delta E$ in the continuum limit by simulating the model at several finite $J=12,...,20$ and $N=30,...,100$. (a) For fixed $N=100,V_{\text{nn}}^{\prime}/\chi=40.53,\lambda_{\kappa}^{\prime}/\chi=0.20$ we perform a linear fit (solid black line) of the energy gap $\Delta E_{\text{latt}}$ (dark green dots, in lattice units) against $1/J(J+1)$, and extrapolate the value in the large spin-length limit $J\rightarrow\infty$ (light green square). (b) The values of the first extrapolation (light green squares) are then fitted against $1/N$ to extrapolate the result in the thermodynamic limit $N\rightarrow\infty$ (light blue plus). The red circle identifies the asymptotic value extrapolated in (a). (c) The data obtained from the fit in $1/N$ corresponds to the continuum limit value $\Delta E_{\text{QFT}}$ (in QFT units) as evidenced by the correct scaling behaviour w.r.t. the bare mass $aM_{0}$. The dashed line represents the theoretical prediction [Eq. (13)] for the mass gap in the simulated QFT, the sine-Gordon model [Eq. (8) for $d=1$]. The red circle identifies the asymptotic value extrapolated in (b). (d) The phase diagram of the sine-Gordon model. The coupling $\beta^{2}$ distinguishes different regimes of the QFT. A BKT transition at $\beta^{2}=8\pi$ separates a gapless ($\beta^{2}>8\pi$) from a gapped region ($\beta^{2}<8\pi$). In the latter the fundamental excitations are (quantum) solitons of mass $M$. Two solitons interact repulsively for $\beta^{2}>4\pi$, while they attract for $\beta^{2}<4\pi$, giving rise to bound states (breathers) with mass $m_{n}$ Eq. (13). The maximal number $n$ of breathers is fixed by $\beta^{2}$ [Eq. (13)]. The mass gap extrapolations from Figs. (a)-(b) correspond to the lowest breather mass $m_{1}$ deep in the attractive regime with $\beta^{2}\approx 0.1\times\pi$, as highlighted by the red circle.

### v2

@@ Abstract
Simulating the real-time dynamics of quantum field theories (QFTs) is one of the most promising applications of quantum simulators. Regularizing a bosonic QFT for quantum simulation purposes typically involves a truncation in Hilbert space in addition to a discretization of space. Here, we discuss how to perform such a regularization of scalar QFTs by explicitly constructing suitable many-body lattice Hamiltonians using multi-level or qudit systems, and show that this enables quantitative predictions in the continuum limit by extrapolating results obtained for large-spin models. With extensive matrix-product state simulations, we numerically demonstrate the sequence of extrapolations that leads to quantitative agreement of observables for the integrable sine-Gordon (sG) QFT. We further show how to prepare static and moving soliton excitations, and analyze their scattering dynamics in the continuum limit, in agreement with a semi-classical model and with quantitative analytical predictions. Finally, we illustrate how a non-integrable perturbation of the sG model gives rise to dynamics reminiscent of string breaking and plasma oscillations in gauge theories. Our methods are directly applicable in state-of-the-art analog quantum simulators, opening the door to quantitatively investigating a wide variety of scalar field theories and tackling long-standing questions in non-equilibrium QFT like the fate of the false vacuum.
@@ I Introduction
Quantum field theories (QFTs) play an essential role in theoretical physics, ranging from effective low-energy descriptions in condensed matter [1] to fundamental models of nature in particle physics [2]. Understanding the physics of QFTs, especially in out-of-equilibrium scenarios, is often guided by computationally challenging numerical simulations, and therefore QFTs constitute an ideal target for quantum simulation [3, 4]. Recent quantum simulation experiments are starting to probe the boundary of classical simulations [5, 6, 7, 8, 9], with the exciting prospect to provide genuine insights into the equilibrium and non-equilibrium properties of quantum many-body systems. However, it remains an outstanding challenge to faithfully simulate physical phenomena of QFTs, such as string-breaking in gauge theories [10, 11, 12, 13].
To clarify the underlying conceptual challenges, recall that QFTs involve continuous variables living on continuous space-time instead of, e.g., discrete spin models on lattice systems. In condensed matter or statistical physics, QFTs often play the role of effective theories, applicable at large distances where the irrelevance of microscopic details and the emergence of a continuum description is explained by the renormalization group. Conversely, space-time lattice discretizations exploit this interpretation of QFTs and provide not only a rigorous non-perturbative definition, but also a computational paradigm that underlies, e.g., lattice QCD [14]. For quantum simulations, which are often set up in a Hamiltonian formulation, these well-established classical techniques need to be fundamentally adapted [15, 16, 17, 18, 19]. An important open problem is to find practically useful truncations of field variables which is unavoidable to represent them on finite-dimensional Hilbert spaces that can be controlled by synthetic quantum systems [20, 21]. How to perform controlled truncations of continuous field variables in order to obtain quantitative results in the continuum limit of QFTs is the central question of this work.
Recent research has mainly addressed this question in the context of universal qubit-based quantum computers [22, 23, 24]. Despite the rapid progress in experimental quantum computing, simulating QFTs in the continuum limit currently remains out of reach of digital approaches due to the large local Hilbert spaces and system sizes required. A promising direction to save resources employs qudit (multi-level) systems [25, 26, 27, 28] as realized with trapped ions [29, 30] or proposed with Rydberg atoms [31, 32, 33]. At the same time, a limited class of field theories is accessible in analog quantum simulators based on continuous quantum gases [34, 35, 36].
@@
Figure 1: Overview. (a) $d$-dimensional square lattice of particles with nearest-neighbor interactions. We assume each particle to be described by a large spin of length $J$, with $\langle\hat{J}_{+}\rangle\sim\langle e^{i\hat{\varphi}}\rangle$ and $\langle\hat{J}_{z}\rangle\sim\langle\hat{\pi}\rangle$ identifying field theory and lattice operators, as described in Sec. II. Continuum physics is extracted by a sequence of extrapolations in spin length $J$, system size NN and lattice spacing $a$ [see Fig. 2]. (b) Sketch of real-time scattering dynamics from large-spin models. We show how to faithfully implement 1. the vacuum (Sec. III), 2. quasi-particle wave-packets and 3. their propagation (Sec. IV), as well as 4. collisions (Sec. V resp. VI), resulting in phase shifts resp. particle production for the example of an integrable resp. non-integrable 1D QFT.
In this work, we lay out a viable path to quantitatively recover the continuum limit of scalar quantum field theories with an explicit truncation of the bosonic field variables. Motivated by a recent proposal to employ large-spin Rydberg atom arrays [33], our approach is based on a family of generalized large-spin Heisenberg lattice Hamiltonians. In a bottom-up approach, we show how appropriate extrapolations utilizing models of increasing spin length [illustrated in Fig. 1(a)] enable the accurate implementation of general scalar field theories, including variants of the paradigmatic sine-Gordon (sG) model. Our results conclusively show that the large-spin regularization provides a clean mapping between microscopic and continuum observables with a precise and systematic understanding how the continuum limit emerges, which enables quantum simulations of QFTs in regimes where no classical benchmarks are available. We emphasize that the conceptual aspects of this regularization are applicable to any quantum simulation that can realize Hamiltonian dynamics, although the specific form of our model Hamiltonian makes it especially well suited for an analog implementation.
@@
To demonstrate our proposal, we perform extensive benchmark simulations employing tensor-network methods for the example of the 1D sG model where exact analytical results are available. In equilibrium, we find quantitative agreement for ground and excited state correlation functions in the continuum limit, validating the correct identification of microscopic degrees of freedom with QFT variables. Out of equilibrium, we discuss how to create solitonic excitations as a feasible alternative to preparing single-particle wave-packets, and find that their spreading induced by quantum fluctuations is well described by a phenomenological semi-classical model. We then simulate the non-equilibrium dynamics of scattering pairs of soliton wave-packets, in agreement with predictions from the known S-matrix. To the best of our knowledge this constitutes one of the first numerically exact simulations of non-equilibrium scattering dynamics in the continuum limit of a bosonic QFT (see also Ref. [37] where scattering in a strongly-coupled Ising field theory is simulated and Ref. [38] which focused on the sG model). As an outlook, we consider a perturbed sG model, where a quantum simulator would provide access to a spatially and temporally resolved picture of scattering dynamics in regimes that potentially lie beyond the reach of traditional methods.
@@
Before delving into the details, the following Section II provides an overview of our approach, including a discussion of its broad applicability and relevance for emulating condensed matter and high-energy physics. We also provide a brief summary of our main results, and the structure of the rest of this paper, which is also illustrated in Fig. 1. Taken together, our results provide a concrete roadmap towards quantitative quantum simulations of field theories.
@@ II Scalar field theories from large-spin models
defined in terms of the continuum field operators $\hat{\varphi}(x)$ and their conjugate momentum $\hat{\pi}(x)$ which fulfill the canonical commutation relation $\hat{\varphi}(x),\hat{\pi}(y)]=i\delta(x-y)$. Different models are distinguished by the form of self-interactions as described by the potential $V$.
@@ II.1 Large-spin lattice models
In App. B, we briefly review a possible experimental implementation of the above Hamiltonian using the SO(4)-symmetric high-dimensional manifold of Rydberg atoms recently discussed in Ref. [33]. In this context the spin length $J=\frac{n-1}{2}$ can be controlled by the principal quantum number $n$ of the employed Rydberg excitation, and all coupling parameters $\Omega,\Delta,\chi,\lambda_{\kappa},V_{\text{nn}}$ can be experimentally tuned by controlling atom-light interactions or rearranging the lattice geometry. Other experimentally relevant setups that give rise to related large-spin models include dipolar molecules [39], magnetic atoms [40, 41] and hybrid digital-analog proposals using qubit clusters of Rydberg atoms [32]. More generally, large-spin models can be implemented with general qudit computers [42] such as those realized with trapped ions [29] as well as hybrid qubit-oscillator processors based on cavity QED [43].
@@ II.2.1 Large-spin limit
[…]
@@
[…]
@@
[…]
Sending $J\rightarrow\infty$ therefore produces the desired lattice field operators. Note that this scaling renders the Ising term ($\propto V_{\text{nn}}\hat{J}_{z}^{(i)}\hat{J}_{z}^{(j)}$) in $\hat{H}_{\text{latt}}$ insignificant in comparison to the “flip-flop” interactions ($\propto V_{\text{nn}}\hat{J}_{+}^{(i)}\hat{J}_{-}^{(j)}$), see App. F.1.
@@ II.2.2 Numerical extrapolations
Conceptually, the large-$J$ limit of the $d$-dimensional lattice Hamiltonian $\hat{H}_{\text{latt}}$ can be translated to a continuum model by increasing the number of lattice sites NN and decreasing a fictitious lattice spacing $a$, while keeping the total volume $V=L^{d}$ with $L=Na$ fixed. Afterwards one can also remove the IR regulator by sending $L\rightarrow\infty$.
@@
To obtain the results shown in Fig. 2, we start by fixing the microscopic lattice parameters of $\hat{H}_{\text{latt}}$ at finite system size $N$, and compute the expectation value of the relevant observables for varying $J$. The numerical results are then fitted with a linear function in $1/J(J+1)$ [see Fig. 2(a)], from which we retrieve the results in the asymptotic limit $1/J(J+1)\rightarrow 0$, equivalent to the large spin-length limit $J\rightarrow\infty$ [see App. D for further details]. In a second step, the asymptotic values obtained for different $N$ are fitted with a quadratic function in $1/N$ [see Fig. 2(b)]. From this additional fit we retrieve the thermodynamic limit $1/N\rightarrow 0$, equivalent to $N\rightarrow\infty$. For small enough lattice spacing $a\rightarrow 0$, we expect scaling behaviour indicative of the continuum limit, which can be compared to theoretical expectations.
@@
Note that we do not tune the lattice spacing $a$ as part of the physical geometry, but we instead vary a dimensionless control parameter $aM_{0}$ – which is a function of the microscopic parameters as discussed below. This parameter sets the mass scale $M_{0}$ of the simulated QFT at the UV cutoff $1/a$ determined by our microscopic lattice model. When all extrapolations are done correctly for sufficiently small values of $aM_{0}$, the system develops a large correlation length $\xi\sim 1/{M_{0}}\gg a$, such that the extrapolated results become insensitive to the lattice regularization, and indeed agree with the continuum predictions. We emphasize that extrapolating $J\rightarrow\infty$ in the first step is an essential ingredient that allows us to perform this analysis. This is in contrast to a fixed small spin length, where the accessible QFT spectrum would be severely truncated to very low energies.
@@
Figure 2: Extrapolating to the continuum limit of the simulated sine-Gordon model. (a)-(c) Extrapolation to the continuum value of the mass gap $\Delta E$. We consider the $d=1$ lattice Hamiltonian $\hat{H}_{\text{latt}}$ Eq. (II.1) and determine the mass gap $\Delta E$ in the continuum limit by simulating the model at several finite J=12,,20J=12,...,20 and N=30,,100N=30,...,100. (a) For fixed $N=100,V_{\text{nn}}^{\prime}/\chi=40.53,\lambda_{\kappa}^{\prime}/\chi=0.20$ we perform a linear fit (solid black line) of the energy gap $\Delta E_{\text{latt}}$ (dark green dots, in lattice units) against $1/J(J+1)$, and extrapolate the value in the large spin-length limit $J\rightarrow\infty$ (light green square). (b) The values of the first extrapolation (light green squares) are then fitted against $1/N$ to extrapolate the result in the thermodynamic limit $N\rightarrow\infty$ (light blue plus). The red circle identifies the asymptotic value extrapolated in (a). (c) The data obtained from the fit in $1/N$ corresponds to the continuum limit value $\Delta E_{\text{QFT}}$ (in QFT units) as evidenced by the correct scaling behaviour w.r.t. the bare mass $aM_{0}$. The dashed line represents the theoretical prediction [Eq. (13)] for the mass gap in the simulated QFT, the sine-Gordon model [Eq. (8) for $d=1$]. The red circle identifies the asymptotic value extrapolated in (b). (d) The phase diagram of the sine-Gordon model. The coupling $\beta^{2}$ distinguishes different regimes of the QFT. A BKT transition at $\beta^{2}=8\pi$ separates a gapless ($\beta^{2}>8\pi$) from a gapped region ($\beta^{2}<8\pi$). In the latter the fundamental excitations are (quantum) solitons of mass $M$. Two solitons interact repulsively for $\beta^{2}>4\pi$, while they attract for $\beta^{2}<4\pi$, giving rise to bound states (breathers) with mass $m_{n}$ Eq. (13). The maximal number $n$ of breathers is fixed by $\beta^{2}$ [Eq. (13)]. The mass gap extrapolations from Figs. (a)-(b) correspond to the lowest breather mass $m_{1}$ deep in the attractive regime with $\beta^{2}\approx 0.1\times\pi$, as highlighted by the red circle.